when 0.1 mole of ammonia is dissolved in sufficient water to make 1 litre of solutionThe solution is found to have a hydroxide ion conc of 1.34 * 10 power -3.The dissociation const of ammonia is.?

Grade:12

5 Answers

Naveen Kumar
9 years ago
Molarity of NH3=0.1mole/1Litre=0.1M
The raction can be given as :
NH3+H2O...........=NH4+....+OH-
0.1M.........................0.........10-7M (due to water)........initially
0.1M-x......................x....... .x+10-7 M..........at equilibrium
Here OH- comes from two sources :one from dissociation of water and the other source is the reaction of NH3 with water.
fom the question,x+10-7 M=10-3..M
so x=~10-3M

writting the dissociation constant, K=[NH4+][OH-]/[NH3]
K=x*x/0.1-x
Shriya Mehrotra
9 years ago
sir i didnt get the ans.
Naveen Kumar
9 years ago
Dear student,
please put the value of [OH-]=1.34*10-3M (Given in the question)
Here I did a mistake in putting the value for hydroxide ion.Then the equation will be
x+10-7M=1.34*10-3..M
and proceed further.


Thank you for your comment
Shriya Mehrotra
9 years ago
thanku sir.got the ans.
Raj kumar
6 years ago
When 0.1 mole of NH3 is dissolved in water to make 1.0 L of soluton, the [OH-] of solution is 1.34×10^-3 M. Calculate Kb for NH3. Please sir,give answer.....???